my sum of my digits is 16 the product of my digits is 64 what number am i?
Answer by stanbon(75887)

my sum of my digits is 16 the product of my digits is 64 what number am i?
---
Assuming the number is a 2 digit number, let it be 10t + u
--------
Equations:
t + u = 16
t*u = 64
--------
Substitute for "t" and solve for "u":
t = 16-u
(16-u)u = 64
16u - u^2 = 64
u^2 - 16u + 64 = 0
-------
Factor:
(u-8)^2 = 0
u = 8
----
Solve for "t":
t + 8 = 16
t = 8
-------
The number is 10t+u = 88
=============
Cheers,
Stan H.